OpenStudy (anonymous):

im hoping this this is a trapoziod. what is the area and preimeter>it's in inches.

OpenStudy (anonymous):

?

OpenStudy (anonymous):



OpenStudy (anonymous):

whats written inside?

OpenStudy (anonymous):

10 ft its the dotted line.

OpenStudy (anonymous):

okay Area = ½ * (a + b) * h

OpenStudy (anonymous):

is this a trapoziod or parallelogram?

OpenStudy (anonymous):

its a trapezoid

OpenStudy (anonymous):

are 110 and 48 the answers?

ganeshie8 (ganeshie8):

trap coz, it doesnt have both pairs of opposite sides equal.

OpenStudy (anonymous):

okay. litke the slanty sides? and top and bottom @ganeshie8

OpenStudy (anonymous):

idk lol @miteshchvm im trying to figure out how to do this in area and permieter

OpenStudy (anonymous):

perimeter can be found by adding all the sides and the formula for the area is given above by me

OpenStudy (anonymous):

yeah you are right @miteshchvm

ganeshie8 (ganeshie8):

exactly, here the opposite slanty sides are equal , each equal to 13. but the opposite base sides are not equal, one is equal to 15, and the other is 7. so its not a parallelogram.

OpenStudy (anonymous):

lol i get it now yay! thanks @ganeshie8

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

Thanks for all yalls help :0 !!!!

OpenStudy (calculusfunctions):

Area of a trapezoid is given by A = (1/2)(h)(a + b) where a and b are the lengths of the two parallel sides and h is the height of the trapezoid. The height is perpendicular to the parallel sides. In your diagram a = 15, b = 7, and h = 10. Thus A = (1/2)(10)(15 + 7) A = (1/2)(10)(22) ∴ A = 110 sq. ft. or 15 840 sq. in. Perimeter = 15 + 13 + 7 + 13 ∴ Perimeter = 48 ft. or 576 in. Note 1 ft. = 12 in. 1 (ft)(ft) = (12 in.)(12 in.) ∴ 1 sq. ft. = 144 sq. in. I hope that was helpful.

OpenStudy (anonymous):

it was thank you.

OpenStudy (calculusfunctions):

Welcome!